OpenStudy (anonymous):

Multiply and simplify: (5 + 2√2) (8 - 3√2) 40 - 5√2 28 + √2 46 + 2√2

Parth (parthkohli):

= 5(8 - 3sqrt2) + 2sqrt2(8 - 3sqrt 2) = (40 - 15sqrt2) + (16sqrt2 - 6) = 28 + sqrt2